Consider the following hypothesis test:


H 0 : μ ≤ 13

H a : μ > 13


A sample of 26 provided a sample mean x̄ = 15 and a sample standard deviation s = 5.32.

a) Compute the value of the test statistic.

b) Use the t distribution table to compute a range for the p-value.

c) At α = 0.01, what is your conclusion?

d) What is the rejection rule using the critical value? What is your conclusion?


1
Expert's answer
2022-01-10T10:00:23-0500

"H_0: \mu ≤ 13 \\

H_1: \mu > 13 \\

n= 26 \\

\bar{x} = 15 \\

s = 5.32"

a) Test-statistic

"t = \frac{\bar{x} - \mu}{s / \sqrt{n}} \\

t = \frac{15-13}{5.32/ \sqrt{26}} = 1.92"

b) Degrees of freedom df = 26-1 = 25

we can find p-value using excel function = T.DIST.RT(1.92,25) = 0.0332

c) α=0.01

p-value > α

Fail to reject the null hypothesis.

Conclusion: "\mu ≤ 13" at 0.01 level of significance.

d) One-tailed test

Reject H0 if "t \u2264 -t_{crit}"

"d.f. = 25 \\

α=0.01 \\

t_{crit} = 2.485 \\

1.92> -2.485"

Fail to reject the null hypothesis.

Conclusion: "\mu ≤ 13" at 0.01 level of significance.